OpenStudy (anonymous):

Complete the following using compound future value (Please use only the following provided Table). (Round your final answers to the nearest dollar amount. Omit the "$" sign in your response.) Time 3 years Principal $2,000 Rate 12% Compounded Semiannually Amount ? Interest ?

OpenStudy (anonymous):

so where is the table? the semi annual interest is 6%. Compound this 6 times A=2000(1.06)^6=2837.04 I=2837.04-2000=837.04

OpenStudy (anonymous):

so do you always break down the percentage rate by the number of periods in a 1 year time span? ie, annually =12%, Semi=6%, and quarterly = 3% ?